OpenStudy (anonymous):

Solve by substitution method: 4x + 3y = 13 x + y = 4

OpenStudy (anonymous):

4x + 3y = 13 x + y = 4 ^ is that it?

OpenStudy (anonymous):

Well if: 4x + 3y = 13 x + y = 4 then we should isolate one of the variables first, in either equation. Lets start with the bottom and isolate (x): x + y = 4 x = 4 - y Now we have (x) in terms of (y), so we can go ahead and plug it in to the first equation: 4x + 3y = 13 4(4-y) + 3y =13 16 - 4y + 3y =13 16 - y = 13 -y = -3 y = 3 Now since we know what (y) is, we can use it to solve for (x): x = 4-y x = 4-3 x = 1 so y = 3 and x = 1